OpenStudy (anonymous):

solve for X (1/3)+(1/5)=(1/x)

OpenStudy (nory):

Start by multiplying both sides by x. What do you get then?

OpenStudy (anonymous):

1/3x+1/5x=x^2 idk

OpenStudy (nory):

x/3 + x/5 = 1

OpenStudy (nory):

You are sort of right...

OpenStudy (nory):

Now to get rid of the fractions, you can multiply both sides by 15. (Does that make sense, why I'm doing that?)

OpenStudy (anonymous):

no

OpenStudy (nory):

If I wanted to get rid of the x/3, I would multiply by 3, because 3 * x/3 = x. Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (nory):

And if I wanted to get rid of the x/5, I would multiply by 5, because 5 * x/5 = x. For the same reason. We want to multiply by 3 _and_ 5, so we just multiply by 15. Does that make sense?

OpenStudy (anonymous):

yeah so what do i do next

OpenStudy (nory):

Well, we have 5x + 3x = 15, after multiplying 15. That makes 8x = 15. Now this is just a simple equation. After solving it, what do we get?

OpenStudy (anonymous):

x=1.85

OpenStudy (nory):

Or x = 15/8. And that's our answer!

OpenStudy (anonymous):

awesome thank you

OpenStudy (nory):

No problem